# Choose Your Model Size: Any Compression by a Single Gradient Descent

### v1

@@ Abstract
The adoption of Foundation Models in resource-constrained environments remains challenging due to their large size and inference costs. A promising way to overcome these limitations is post-training compression, which aims to balance reduced model size against performance degradation. This work presents Any Compression via Iterative Pruning (ACIP), a novel algorithmic approach to determine a compression-performance trade-off from a single stochastic gradient descent run. To ensure parameter efficiency, we use an SVD-reparametrization of linear layers and iteratively prune their singular values with a sparsity-inducing penalty. The resulting pruning order gives rise to a global parameter ranking that allows us to materialize models of any target size. Importantly, the compressed models exhibit strong predictive downstream performance without the need for costly fine-tuning. We evaluate ACIP on a large selection of open-weight LLMs and tasks, and demonstrate state-of-the-art results compared to existing factorisation-based compression methods. We also show that ACIP seamlessly complements common quantization-based compression techniques.
Model Compression, Weight Factorization, SVD, Structured Pruning, LLM, Optimization path
@@ 1 Introduction
Figure 1: Compression-performance trade-off curves generated by ACIP on C4 for a variety of LLMs. Each curve is obtained by a single gradient descent run and all compressed models can be materialized without recomputation.
Figure 2: A visual overview of ACIP. 1 The linear layers of the base model are reparametrized in terms of their singular value decomposition $\mathbf{U}\mathbf{M}\boldsymbol{\Sigma}\mathbf{V}^{\top}$, with a (binary) singular value mask $\mathbf{M}$, and a low-rank adapter $\boldsymbol{\Delta}$. 2 The objective function is optimized over the mask parameters $\mathbf{p}$ and adapters $\boldsymbol{\Delta}$, where sparsity is induced on $\mathbf{p}$ by an $\ell_{1}$-penality leading to pruned entries in the mask $\mathbf{M}$. The optimization path of $\mathbf{p}$ gives rise to a score map that determines the global importance of the singular values across the full model. Potential compression errors are compensated by $\boldsymbol{\Delta}$. 3 Based on the parameter scores, the base model can be flexibly compressed to any target size by masking the entries of $\boldsymbol{\Sigma}$. The learned adapters $\boldsymbol{\Delta}$ are used as correction for any compression level.
@@
Post-training compression of Foundation Models, especially large language models (LLMs), promises access to powerful tools where resources are limited, e.g., in e-commerce, mobile deployments, or on the shop floor. Typical reasons for scarcity include constrained access to hardware, monetary limitations, high inference speed requirements, and environmental concerns (Hohman et al., 2024). However, model compression is not for free — it typically comes with a trade-off between model size and desired downstream performance. In order to support decision-making, it is of fundamental importance to find the right balance between these two objectives. In fact, when deploying large-scale models, questions like the following ones arise:
  • I am willing to give up on $X\%$ of model performance, how much can I compress model $Y$?
    My infrastructure allows me to run an $N$ parameter model, what is the best model within this budget?
    What are the minimal hardware requirements for an LLM with a benchmark performance of at least $Z$?
To make model compression more accessible, the algorithmic characterization of the underlying trade-off should be a high priority in the development of new algorithms (Boggust et al., 2025). In the scientific literature, however, this aspect is often secondary, and instead, performance is compared at preset compression levels (Zhu et al., 2024).
Apart from resource-intensive model distillation (Hinton et al., 2015), there are currently two dominating approaches to efficient post-training model compression: (i) model quantization, which limits weight fidelity by reducing numerical precision, and (ii) (un-)structured weight pruning which aims to delete unnecessary or redundant weights from a model (Zhu et al., 2024). While many of these approaches lead to good compression results, they typically allow only for constant compression factors in case of quantization or require specialized hardware, e.g., to implement $n{:}m$-sparsity in unstructured pruning or low-bit representations in quantization. Due to these structural constraints, such methods do not enable a full characterization of the compression-performance trade-off for a given base model.
Inspired by the effectiveness of low-rank adapters in LLMs (Hu et al., 2022), factorization-based compression can be applied without any integral changes to a network’s architecture, by simply reparametrizing linear layers. This particular form of structured pruning allows, in principle, for compression factors on a fine-grained scale without additional hardware constraints. However, the performance of existing algorithms falls behind the above-described compression techniques, so that competitive performance for stronger size-reduction is only achieved with additional fine-tuning (Wang et al., 2024; Yuan et al., 2024; Jaiswal et al., 2024).
In this work, we follow the promising idea of factorization-based compression and find a positive answer to our initial quest:
  • Design an efficient compression procedure that algorithmically determines the compression-performance trade-off for a given pre-trained model.
The basic idea of our proposed method, Any Compression via Iterative Pruning (ACIP), [^1] Pronounced like “a sip” of coffee. is to induce sparsity in the singular values of the weight matrices, which leads to low-rank and therefore compressed parametrizations. Instead of layer-wise updates, we rely on a simple end-to-end objective function that is optimized through stochastic gradient descent. Crucially, the optimization trajectory of the pruned singular values gives rise to a score map representing a global importance ranking across all linear network layers. This allows us to materialize compressed models of any target size without additional computation and produce trade-off curves like in Figure 1. See also Figure 2 for a visualization of ACIP and Section 2 for a more detailed description.
We will demonstrate that ACIP satisfies the following practical requirements, which at the same time, also form the key contributions of our work:
Easy-to-Use & Fast
ACIP allows practitioners to compress pre-trained models to any size from a single optimization run. It is parameter-efficient, requires no hyperparameter tuning and no feature engineering due to its end-to-end nature.
Flexible & Complementary
ACIP can be applied to any model that has linear layers as central building blocks without further structural assumptions. It can be used with arbitrary task loss functions and seamlessly complements quantization-based compression approaches.
Effective & Robust
ACIP exhibits state-of-the-art results among factorization-based pruning approaches. It provides consistent and robust compression-performance trade-offs for a variety of popular LLMs and zero-shot evaluation tasks.
Our main results supporting these claims are presented in Section 4. The interested reader is also referred to the supplementary material, where we address some design decisions and conceptual aspects (Appendix A), provide implementation and runtime details (Appendix B), as well as additional experiments (Appendix C and Appendix D).
@@ 2 Method
ACIP is an easy-to-use method to build models at any compression rate based on an SVD-weight factorization (Section 2.1.2). We leverage ideas from sparse recovery (Section 2.1.3) to iteratively prune model parameters and then use the resulting pruning order to derive a global importance ranking (Section 2.1.1). With these preliminaries at hand, the algorithmic details of ACIP are formulated in Section 2.2.
@@ 2.1.1 Score-based Parameter Pruning
As motivated in the introduction, the goal of this work is to determine the compression-performance trade-off of a pre-trained model. To do this, we use score-based parameter pruning, a framework that has been successfully applied to model compression since the 1980s (LeCun et al., 1989; Hassibi et al., 1993). In score-based pruning, a score map $\boldsymbol{\rho}$ is created that assigns an importance score $\rho_{i}$ to all target parameters $\theta_{i}$. Naturally, this approach gives rise to a (global) ranking of parameters that allows for model compression at any desired rate.
There are many ways to design useful score maps. For example, they can be derived based on curvature at a local optimum (LeCun et al., 1989; Hassibi et al., 1993) or by using hand-designed local features (Sun et al., 2024; Frantar & Alistarh, 2023). In Section 2.2.2, we present a novel, data-driven approach to score maps that does not require any handcrafting or feature engineering.
2.1.2 Compressing Linear Layers
This work builds on the compression of the weight matrices in linear layers of a model. These weights typically account for the largest portion of parameters in deep learning models. We consider linear layers of the form
@@ 2.1.2 Compressing Linear Layers
where $\mathbf{W}$ is a $m\times n$ matrix and $\mathbf{b}$ is a bias term. We aim for a low-rank compression based on a matrix factorization such that
@@
with $\mathbf{P}$ and $\mathbf{Q}$ being matrices of sizes $m\times k$ and $k\times n$, respectively, where $k\ll\min(m,n)$. The layer parametrization in (2) may be interpreted as a (lossy) compression of the parametrization in (1), leading to a smaller memory footprint whenever $k(m+n)<mn$. Note that matrix factorization can be used to compress the parameters of any linear layer, including dense layers as well as efficiently parametrized layers such as convolutional layers (Idelbayev & Carreira-Perpiñán, 2020). In this work, we focus on the compression of dense layers in transformers (Vaswani et al., 2017).
Singular Value Decomposition
For low-rank factorizations in ACIP, we choose the singular value decomposition (SVD) which has been used extensively in model compression (Idelbayev & Carreira-Perpiñán, 2020; Yuan et al., 2024; Wang et al., 2024). SVD factorizes any $m\times n$ weight matrix $\mathbf{W}_{l}$ of rank $r\leq\min(m,n)$ at layer $l$ as
@@
where $\mathbf{U}_{l}$ and $\mathbf{V}_{l}$ are $m\times r$ and $n\times r$ matrices of (orthonormal) singular vectors, and $\boldsymbol{\Sigma}_{l}$ is a $r\times r$ diagonal matrix containing the singular values $\mathbf{s}_{l}^{(i)}>0$. Note that we have used the compact SVD which ignores the null-space vectors of the full SVD. We may reduce the inner dimension $r$ of the SVD to $k<r$ with
$\mathbf{W}_{l}\approx\tilde{\mathbf{U}}_{l}\tilde{\boldsymbol{\Sigma}}_{l}\tilde{\mathbf{V}}_{l}^{\top}$ (4)
by defining $\tilde{\mathbf{U}}_{l}$ and $\tilde{\mathbf{V}}_{l}^{\top}$ to only contain $k$ singular vectors associated with a selected subset of singular values. To preserve the matrix norm of $\mathbf{W}_{l}$, selecting the $k$ largest singular values leads to an optimal approximation of $\mathbf{W}_{l}$ under rank constraints (Mirsky, 1960). It has been previously argued, however, that this approach does not yield satisfactory results in deep learning as is does not take into account the (training) data (Hsu et al., 2022). Different approaches have been presented to address this problem (Hsu et al., 2022; Yuan et al., 2024; Wang et al., 2024; Chen et al., 2021).
@@ 2.1.3 $\ell_{1}$-regularization
Consider a generic loss functional $\mathcal{L}(\mathbf{X};\boldsymbol{\theta})$ evaluated on some model that is specified through a parameter vector $\boldsymbol{\theta}$ and data $\mathbf{X}$. During model training, sparsity in $\boldsymbol{\theta}$ is typically encouraged by solving the penalized optimization problem
@@
which is known as $\ell_{1}$-regularization, or least absolute shrinkage and selection operator (LASSO) in case of (generalized) linear models (Tibshirani, 1996; Hastie et al., 2015). Feature selection plays a key role in this optimization problem. Consider the solution to (5) as a function of regularization parameter $\lambda$: Increasing $\lambda$ leads to a sparser solution in $\boldsymbol{\theta}$. This effectively gives a ranking of feature importances for the solution of (5) through the so-called regularization path (Tibshirani, 1996; Efron et al., 2004; Mairal & Yu, 2012) — an idea that inspired our approach. Furthermore, previous results indicate that optimization paths of iterative schemes can be related to regularization paths (Suggala et al., 2018). Intuitively, this suggests that general optimization paths of iterative schemes for $\ell_{1}$-objectives reveal information about feature importance in the context of sparse models as well.
In this work, we will successively increase $\lambda$ to introduce a higher degree of model compression in terms of parameter sparsity in a controlled manner (see also Remark 2.1 below).

### v2

# Choose Your Model Size: Any Compression of Large Language Models Without Re-Computation

@@ Abstract
The adoption of Foundation Models in resource-constrained environments remains challenging due to their large size and inference costs. A promising way to overcome these limitations is post-training compression, which aims to balance reduced model size against performance degradation. This work presents Any Compression via Iterative Pruning (ACIP), a novel algorithmic approach to determine a compression-performance trade-off from a single stochastic gradient descent run. To achieve parameter efficiency, we use an SVD-reparametrization of linear layers and iteratively prune their singular values with a sparsity-inducing penalty. Importantly, the pruning order of the parameters is used to derive a global score map that allows compressing a model to any target size without re-computation. We evaluate ACIP on a large selection of open-weight LLMs and downstream tasks, demonstrating state-of-the-art results compared to existing factorization-based compression methods. We also show that ACIP seamlessly complements common quantization-based compression techniques.
@@ 1 Introduction
Post-training compression of Foundation Models, especially Large Language Models (LLMs), promises access to powerful tools where resources are limited, e.g., in automotive systems, mobile deployments, or on shop floors Gholami et al. (2022). Typical reasons for resource scarcity include constrained access to hardware, monetary limitations, high inference speed requirements, and environmental concerns Hohman et al. (2024).
The original promise of model compression was to eliminate redundant parameters, resulting in almost lossless methods Han et al. (2016). While working well for models trained on smaller datasets, this hypothesis does not hold up anymore in the era of LLMs and scaling laws Allen-Zhu & Li (2024). For modern “densely trained” models, compression is almost always lossy, leading to a fundamental trade-off between model size and downstream performance. While characterizing this trade-off supports practitioners in deployment decisions Boggust et al. (2025), the scientific literature typically focuses on benchmarks at preset compression levels Zhu et al. (2024). This gap between research and practice implies that, for model users, the process is often perceived as a “black box”, requiring significant expertise and trial-and-error to identify an acceptable setup. We argue for the opposite approach, one that empowers users to seamlessly customize a compression algorithm for their specific use cases.
Any Compression
(a) Conventional Model Compression
(b) Any Compression
Figure 1: Compared to conventional compression algorithms (a), an Any Compression algorithm (b) swaps the computational calibration step and the decision step, so that models of different target sizes can be materialized without re-computation.
With this motivation in mind, we advocate for methods that permit Any Compression of pre-trained models. Here, ‘Any’ signifies an algorithm’s ability to scale a given base model to an arbitrary target size, guided by the user’s specific needs and limitations, rather than the algorithm dictating possible sizes. To facilitate decision-making, such an algorithm must efficiently reveal the compression-performance trade-off without extensive re-computation. In existing post-training compression approaches, we identify two practical challenges to achieving Any Compression.
Problem 1 (Preset compression rates).
The size reduction of large Foundation Models is a prominent research area, with existing methods largely falling into three categories: knowledge distillation (training a smaller student model), quantization (reducing numerical precision), and parameter pruning (removing redundant weights) — for a comprehensive survey, we refer to Zhu et al. (2024). While effective, these approaches often impose constraints that conflict with the goal of achieving Any Compression as they are restricted to preset (discrete) compression factors. Indeed, knowledge distillation is limited to a single compression rate defined by the fixed size of the student model Hinton et al. (2015). Quantization is limited by fixed bit-length reductions (typically from 16-bit to 4-bit or 8-bit representations) Gholami et al. (2022); Dettmers et al. (2023). Similarly, (unstructured) parameter pruning often relies on rigid sparsity patterns (e.g., n:mn{:}m sparsity) to ensure efficient memory allocation and hardware acceleration, thereby restricting the possible model sizes Choquette et al. (2021). In practice, preset compression rates are especially problematic when deploying under a specific, non-discrete constraint, e.g., a fixed memory budget, forcing users into a costly “guess-and-check” cycle.
Solution (Structured pruning by weight factorization).
Weight factorization is a technique where (linear) model weights are decomposed into several sub-matrices Eckart & Young (1936). This approach enables fine-grained model compression and high parameter efficiency by pruning the inner dimensions of the resulting matrices Yuan et al. (2024); Wang et al. (2024). A key advantage is its compatibility with standard hardware, since it only requires basic matrix-vector multiplications.
Problem 2 (Different compression rates require re-computation).
Most existing post-training compression techniques are inherently inefficient for exploring the trade-off between model size and performance. Conventionally, a user selects a target compression rate and other hyperparameters, initiating a costly computational step that can take minutes to hours Frantar et al. (2023); Sun et al. (2024); Yuan et al. (2024). Each new compression rate requires repeating this entire process, making a comprehensive evaluation of the trade-off landscape impractical (see Figure˜1(a)).
Solution (Any Compression through amortization).
We argue that a reversed workflow is preferable: a single, upfront computational investment that enables the subsequent materialization of a model at any compression rate in almost real-time (see Figure˜1(b)). The initial effort, which could be handled by the model supplier, would empower users to efficiently select an optimal model instance for their needs at a negligible cost. The challenge, therefore, is to design algorithms that are compatible with this “compute once, compress dynamically” approach.
Figure 2: A visual overview of ACIP. 1 The linear layers of the base model are reparametrized in terms of their singular value decomposition $\mathbf{U}\mathbf{M}\boldsymbol{\Sigma}\mathbf{V}^{\top}$, with a (binary) singular value mask $\mathbf{M}=\mathbf{M}(\mathbf{p})$ and a low-rank adapter $\boldsymbol{\Delta}$. 2 An objective function is optimized via gradient descent over the mask parameters $\mathbf{p}$ and adapters $\boldsymbol{\Delta}$, where sparsity is induced on $\mathbf{p}$ by an increasing $\ell_{1}$-penalty. This leads to pruned entries in the mask $\mathbf{M}(\mathbf{p})$. The optimization path of $\mathbf{p}$ gives rise to a score map that determines the global importance of the singular values across the full model. Potential compression errors are compensated by $\boldsymbol{\Delta}$. 3 Based on the parameter scores, the base model can be flexibly compressed to any target size by masking the entries of $\boldsymbol{\Sigma}$. The learned adapters $\boldsymbol{\Delta}$ are used as correction for any compression level.
Contributions and Overview
In this work, we introduce Any Compression via Iterative Pruning (ACIP), [^1] Pronounced like ‘a sip’ of coffee. which is specifically developed to address the above problems. To the best of our knowledge, ACIP is the first algorithm that enables large-scale model compression to any size in real-time without requiring re-computation or re-calibration.
To overcome ˜1, ACIP follows a low-rank factorization strategy, based on pruning singular values of large linear layers. This particular form of structured parameter pruning allows for fine-grained compression levels and is parameter-efficient at the same time, as only singular values are altered. Compared to quantization and conventional (unstructured) pruning, an efficient implementation does not come with any restrictions to the underlying hardware, since it is purely based on standard matrix-vector multiplications.
ACIP solves ˜2 by explicitly decoupling an (optimization-based) pruning stage from the actual compression stage. The former can be viewed as a data-dependent calibration step, which estimates the global importance of all target parameters (the singular values of the base model layers) through an iterative pruning scheme. The resulting score map is then used to implement a simple compression step that enables the instantiation of models of any desired size without further computational costs. The detailed methodology and technicalities of ACIP are presented in Section˜2 (see Figure˜2 for a visual overview).
We empirically demonstrate the effectiveness of ACIP on a range of recent LLMs in Section˜3, accompanied by a series of additional experiments in the supplementary material (Appendices˜C and D). In particular, we verify that our approach outperforms other factorization-based methods on multiple benchmarks (Section˜3.2) and seamlessly complements quantization-based compression techniques (Section˜3.4). In the spirit of scaling laws Kaplan et al. (2020), ACIP provides consistent and robust size-performance trade-offs, which allows model users to predict downstream capabilities from a few data points.
Finally, we put our results in the broader context of related literature in Section˜4 and conclude in Section˜5 by outlining limitations as well as promising avenues for future research that build on our contributions.
@@ 2 Method
Figure˜2 provides a schematic overview of Any Compression via Iterative Pruning (ACIP). In its initial stage, ACIP builds a reparametrization of large linear network layers by a singular value decomposition (SVD), which enables model compression through rank reduction (see Section˜2.2.1 for details). Unlike existing SVD-based methods Idelbayev & Carreira-Perpiñán (2020); Yuan et al. (2024); Wang et al. (2024), Any Compression is achieved by decoupling the pruning and compression stages (cf. Figure˜1(b)). More specifically, we construct a score map that establishes a global importance ranking of singular values across all linear layers within the network. The score map is derived by running a (stochastic) gradient descent on a sparsity-inducing objective, using the pruning order of the singular values as a proxy for feature importance (see Section˜2.2.2). In an independent step, the score map can then be used to compress the base model to any desired size without re-computation (see Section˜2.2.3).
@@ 2.1 Preliminaries
This section provides several preliminaries to better understand the algorithmic details of ACIP, which are presented in Section˜2.2.
@@ 2.1.1 Score-Based Parameter Pruning
As motivated above, the overarching goal of this work is to allow Any Compression of a pre-trained model by decoupling the computational stage from the compression stage. To this end, we use score-based parameter pruning, a framework that has been successfully applied to model compression since the 1980s LeCun et al. (1989); Hassibi et al. (1993). In score-based pruning, a score map $\boldsymbol{\rho}$ is created that assigns an importance score $\rho_{i}$ to each target parameter $\theta_{i}$. Naturally, this approach gives rise to a (global) ranking of parameters that allows for model compression at any desired rate.
There are many ways to design useful score maps. For example, they can be derived based on curvature at a local optimum LeCun et al. (1989); Hassibi et al. (1993) or by using hand-designed local features Sun et al. (2024); Frantar & Alistarh (2023). ACIP takes a novel, data-driven approach to score maps that does not require any handcrafting or feature engineering (see Section˜2.2.2).
2.1.2 Low-Rank Compression of Linear Layers
Linear layers are the molecular building blocks of modern machine learning models, typically accounting for more than $90\%$ of model parameters in transformers (Vaswani et al., 2017), which makes them a natural target for compression. Accordingly, we consider linear layers of the form
@@ 2.1.2 Low-Rank Compression of Linear Layers
where $\mathbf{W}$ is an $m\times n$ matrix, $\mathbf{b}$ is a bias term, and $\mathbf{x}$ and $\mathbf{y}$ are layer inputs and outputs, respectively. In this work, we specifically aim for a low-rank compression based on a matrix factorization such that
@@
where $\mathbf{P}$ and $\mathbf{Q}$ are matrices of sizes $m\times k$ and $k\times n$, respectively, and $k\ll\min(m,n)$. The layer parametrization in (2) may be interpreted as a (lossy) compression of the parametrization in (1), leading to a smaller memory footprint whenever $k(m+n)<mn$. Note that matrix factorization can be used to compress the parameters of any linear layer, including dense layers as well as efficiently parametrized layers such as convolutional layers Idelbayev & Carreira-Perpiñán (2020).
Singular value decomposition.
To determine suitable low-rank factors $\mathbf{P}$ and $\mathbf{Q}$ in (2), we follow recent work on (large) model compression Idelbayev & Carreira-Perpiñán (2020); Yuan et al. (2024); Wang et al. (2024) and leverage a singular value decomposition (SVD). SVD factorizes an $m\times n$ weight matrix $\mathbf{W}_{l}$ of rank $r\leq\min(m,n)$ at layer $l$ as
@@
where $\mathbf{U}_{l}$ and $\mathbf{V}_{l}$ are $m\times r$ and $n\times r$ matrices of (orthonormal) singular vectors, and $\boldsymbol{\Sigma}_{l}$ is a $r\times r$ diagonal matrix containing the singular values $\mathbf{s}_{l}^{(i)}>0$. Note that we consider the compact SVD here, which ignores the null-space vectors of the full SVD. We may reduce the inner dimension $r$ of the SVD to $k<r$ with
$\mathbf{W}_{l}\approx\tilde{\mathbf{U}}_{l}\tilde{\boldsymbol{\Sigma}}_{l}\tilde{\mathbf{V}}_{l}^{\top},$ (4)
by defining $\tilde{\mathbf{U}}_{l}$ and $\tilde{\mathbf{V}}_{l}^{\top}$ to only contain $k$ singular vectors associated with a selected subset of singular values. Measuring the approximation error in terms of (Frobenius) matrix norm, selecting the $k$ largest singular values leads to an optimal approximation of $\mathbf{W}_{l}$ under rank constraints (Mirsky, 1960). It has been previously argued, however, that this approach does not yield satisfactory results in deep learning as is does not take into account the underlying (training) data and downstream task Hsu et al. (2022). Different approaches have been presented to address this problem (Hsu et al., 2022; Yuan et al., 2024; Wang et al., 2024; Chen et al., 2021).
@@ 2.1.3 $\ell_{1}$-Regularization
Consider a generic loss functional $\mathcal{L}(\mathbf{X};\boldsymbol{\theta})$ evaluated on some model that is specified by a parameter vector $\boldsymbol{\theta}$ and data $\mathbf{X}$. During model training, sparsity in $\boldsymbol{\theta}$ is typically encouraged by solving the penalized optimization problem
@@
which is known as $\ell_{1}$-regularization, or least absolute shrinkage and selection operator (LASSO) in case of (generalized) linear models Tibshirani (1996); Hastie et al. (2015). Feature selection plays a key role in this optimization problem. Indeed, increasing $\lambda$ in (5) leads to sparser solutions $\hat{\boldsymbol{\theta}}(\lambda)$. This effectively gives rise to a feature importance ranking for the solution of (5) through the so-called regularization path Tibshirani (1996); Efron et al. (2004); Mairal & Yu (2012) — an idea that inspired the approach of ACIP. Here, ‘feature importance’ is understood as the relative contribution to the (task- and data-specific) loss $\mathcal{L}\bigl(\mathbf{X};\boldsymbol{\theta}\bigr)$, i.e., a feature is considered more important if its removal causes a larger increase in the loss. In particular, different choices of $\mathcal{L}$ may lead to different solution paths and feature rankings. Furthermore, previous results indicate that optimization paths of iterative schemes can be related to regularization paths (Suggala et al., 2018). Intuitively, this suggests that general optimization paths of iterative schemes for $\ell_{1}$-objectives reveal information about feature importance in the context of sparse models as well.
In ACIP, we will successively increase $\lambda$ to introduce a higher degree of model compression in terms of parameter sparsity in a controlled manner (see also Remark˜2.1 below).